OpenStudy (anonymous):

Solve. 14x2 + 42x + 70 = 0

OpenStudy (anonymous):

the best way is to reduce them first before you use the quadratic formula,

OpenStudy (anonymous):

14x^2 + 42x + 70 = 0 divide by 14 x^2 + 3x + 5 = 0 (x+3/2)^2+5-9/4=0 (x+3/2)^2+11/4=0 (x+3/2)^2=-11/4 (x+3/2)=-+sqrt(-11/4) x=-+sqrt(-11/4)-3/2 x=-1.5+sqrt(11/4)i=-1.5+1.65i x=-1.5-sqrt(11/4)i=-1.5-1.65i